Please help me solve this problem:
A father and son are 50 years old and 20 years old, respectively.
Find the ratio of the the father's age to the son's age, when:
a. Then the son was 10 years old? How old was the father?
b. The the father was 35 years old? How old was the son?
c. The son will be 40 years old? How old will be the father?
d. The father will be 75 years old? How old will be the son?
Thank you very much.

Answer by rothauserc(4718)

the ratio of the father's age to the son's age = 50/20 = 5/2
:
a) (5/2) * 10 = 25, the father is 25 years old
:
b) (5/2) * x = 35
x = 70/5 = 14, the son is 14 years old
:
c) (5/2) * 40 = 100, father will be 100 years old
:
d) (5/2) * x = 75
x = 150/5 = 30, son is 30 years old
: